OpenStudy (anonymous):

-(27/64y to the sixth power) to the 2/3 power

OpenStudy (blurbendy):

(-(27/64 y)^6)^2/3 ??

OpenStudy (anonymous):

only one set of parentheses

OpenStudy (blurbendy):

-(26/64y)^6 <--- then ALL of that to the 2/3 power, or ^6^2/3

OpenStudy (anonymous):

all of that to the 2/3 power its 27 not 26

OpenStudy (blurbendy):

tomatoes, tamatoes, hold on.

OpenStudy (blurbendy):

you want to solve for y, or just simplify?

OpenStudy (anonymous):

i keep getting it wrong

OpenStudy (anonymous):

simplify

OpenStudy (blurbendy):

Well you can rewrite this as \[1/3(-(27y/64)^6)^2\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

only 64y has the sixth power by the way. sorry

OpenStudy (blurbendy):

\[-(27/64y^6)^2/3\]

OpenStudy (anonymous):

the question is:Simplify the expression. The variable represents a positive real number.

OpenStudy (blurbendy):

^2/3

OpenStudy (blurbendy):

like that?

OpenStudy (anonymous):

yup

OpenStudy (blurbendy):

then you can rewrite that as -1/3((27)/(64y^6))^2 = -243 / 4096y^12

OpenStudy (anonymous):

that was wrong, i think they are looking for a radical

OpenStudy (blurbendy):

bump your question, because that's how i would have simplified it.

OpenStudy (blurbendy):

okay, maybe this what they mean if they are looking for a radical -(27/64y^6)^3 = \[- \sqrt[3]{729 / 4096y^(12)} = -9/4096y^4\]19683 / 262144y^10

OpenStudy (blurbendy):

oops, ignore the last part with 19683

OpenStudy (blurbendy):

by the way, are you sure you entered the first answer correctly? because that should be right.